OpenStudy (anonymous):

8(7x - 3) = -6 Answer As A Fraction

OpenStudy (anonymous):

8(7x - 3) = -6 56x - 24 = -6 56x = 24 -6 56x = 18

OpenStudy (anonymous):

Thank U so much now i will get into harverd

OpenStudy (anonymous):

x = 18/56

OpenStudy (anonymous):

you are very welcome :)

OpenStudy (anonymous):

x = 18/56 = 9/28